### v3

# A Comparison of Relativistic Coupled Cluster and Equation of Motion Coupled Cluster Quadratic Response Theory

Abstract
We present the implementation of relativistic coupled cluster quadratic response theory (QR-CC), following our development of relativistic equation of motion coupled cluster quadratic response theory (QR-EOMCC) [X. Yuan et al., J. Chem. Theory Comput. 2023, 19, 9248–9259]. These codes, which can be used in combination with relativistic (2- and 4-component based) as well as non-relativistic Hamiltonians, are capable of treating both static and dynamic perturbations for electric and magnetic operators. We have employed this new implementation to revisit the calculation of static and frequency-dependent first hyperpolarizabilities of hydrogen halides (HX, X=F-Ts) and the Verdet constant of heavy noble gas atoms (Xe, Rn, Og) and of selected hydrogen halides (HF to HI), in order to investigate the differences and similarities of QR-CC and the more approximate QR-EOMCC. Furthermore, we have determined the relative importance of scalar relativistic effects and spin-orbit coupling to these properties, through a comparison of different Hamiltonians, and extended our calculations to superheavy element species (HTs for hyperpolarizabilities, Og for the Verdet constant). Our results show that as one moves towards the bottom of the periodic table, QR-EOMCC can yield rather different results (hyperpolarizabilities) or perform rather similarly (Verdet constant) to QR-CC. These results underscore the importance of further characterizing the performance of QR-EOMCC for heavy element systems.
keywords:
[…]
1 Introduction
External perturbations such as those induced by electric or magnetic fields allow us to probe the electronic structure of atoms, molecules and materials, with different perturbations providing us with complementary information about the system norman_principles_2018, Jaszuski2015. The first-order (linear) responses to such a perturbation are particularly sought-after quantities, since from them one can obtain information about electronically or vibrationally bishop_molecular_1990 excited states though processes involving a single photon, or magnetic properties such as NMR shieldings or indirect spin-spin coupling constants helgaker1999ab, vaara2007theory, helgaker2008quantum to name but a few.
Properties arising from the interaction with two or more photons papadopoulos_non-linear_2006, he2008multiphoton at a time are also of great interest be it for technological applications in non-linear optics Norman2006, boyd2020nonlinear or for speciation through two-photon absorption cronstrand_multi-photon_2005, as the differences in selection rules compared to one-photon processes allow for probing different states.
Theoretical models are indispensable to interpret experimental results, and the last fifty years have seen the development helgaker_recent_2012, helgaker_molecular_2014 of approaches based on ab initio molecular electronic structure theory gain popularity in applications, thanks in no small part to the development of response theory norman_perspective_2011, Pedersen2015, Jaszuski2015 and its application to different models such as multi-configurational self-consistent field theory dalgaard1980time, olsen_linear_1985, Hettema1992, density functional theory Casida1996, and coupled cluster theory Monkhorst2009, dalgaard_aspects_1983, Koch1990a, Koch1990b, Koch1994, kobayashi1994calculation, kondo_orthogonally_1995, piecuch_property_1995. The use of the time-averaged quasienergy formalism langhoff1972aspects, rice1991calculation, sasagane1993higher, Koch1994, Christiansen1995a, Christiansen1995b, Christiansen1996, Hattig1997, coriani_coupled_1997, Christiansen1998, Gauss1998, kauczor2013communication, khani_uv_2019, hattig_cc2_2000 has provided not only a unified framework, with which to derive response for both variational and non-variational approaches, but also a way to enable the development of embedding approaches in which variational and non-variational electronic structure methods are combined Hofener2012, Hofener2012b, Hofener2013, Hofener2016, Niemeyer2025, Myhre2014, Myhre2016, Goletto2021 to calculate molecular properties in large systems within a fully quantum mechanical description in a cost-effective manner.
In the case of coupled cluster theory, one can identify two main approaches, one which employs an exponential parametrization for the time evolution of the cluster operators (CC response)Monkhorst2009, dalgaard_aspects_1983, Koch1990a, Koch1990b, Koch1994, kobayashi1994calculation, Christiansen1995a, Christiansen1995b, Christiansen1996, Hattig1997, coriani_coupled_1997, Christiansen1998, kondo_orthogonally_1995, piecuch_property_1995, spirko_molecular_1996, piecuch_molecular_1996, piecuch_convergence_1997, and another based on the equation of motion formalism (EOMCC response)Stanton1993, Stanton1993b, Sekino1995, gauss_coupledcluster_1995, Rozyczko1997, Sekino1999, Crawford2009, Nanda2015, nanda_static_2016, nanda_communication_2018 for which a linear parametrization is employed to obtain molecular properties.
From the outset, key differences between the two approaches have been identified with respect to separability Koch1994, Stanton1994, caricato2009difference, with as consequence that CC transition moments are size-intensive whereas EOMCC ones are not. More recently the interconnections between these two approaches were further investigated from a formal point of view pawlowski_molecular_2015, coriani_molecular_2016.
In spite of such differences, the calculation of response properties within the EOMCC formalism remains appealing from both the implementation and the application points of view faber2018resonant, andersen2022cherry, andersen2022probing, as simplifications in the equations also translate into a somewhat more modest computational cost. This calls for an understanding of how the two approaches differ in practical calculations. In the literature comparisons between linear response CC (LR-CC) and EOMCC (LR-EOMCC) properties Sekino1994, Sekino1999, Crawford2009, coriani_molecular_2016 so far suggest that, in spite of the lack of size-extensivity of EOM-CC transition moments, in practice the two approaches yield quite similar results.
Similar conclusions have been reached when comparing quadratic response CC (QR-CC) and EOMCC (QR-EOMCC) results for dipole hyperpolarizabilties Rozyczko1997, Sekino1995, coriani_molecular_2016, though comparisons were initially limited to a few systems. More recently, an extensive benchmark study on excited state transition moments irek2025, which also probe quadratic response functions, seem to confirm the tendencies observed previously. These previous evaluations were carried out for valence properties of molecules containing light elements, that is, occupying the first two rows of the periodic table. For properly representing the molecular properties of heavier systems it is necessary to account for relativistic effectsPyykko1988a, saue_relativistic_2011, Pyykk2012, Autschbach2012a, liu2017handbook, Schwerdtfeger2020, Smits2023 and to the best of our knowledge there is currently no data in the literature assessing the differences between the two approaches beyond the second row of the periodic table.
Response theory can be used in combination with relativistic HamiltoniansSaue2002_PostDirac, and implementations for mean-field approaches (Hartree-Fock and DFT) are for example available for the calculation of frequency-dependent linear and non-linear response properties for 4-component-based Hamiltoniansvisscher_4-component_1997, Saue:Jensen:JCP2003, norman_quadratic_2004, salek2005linear, Henriksson:2005, Henriksson_JCP2006, Tellgren_JCP2007, bast:JCP2009, Bast_ijqc2009, Bast2009b, Villaume_JCP2010. To go beyond mean-field, relativistic coupled cluster theory Liu2021, visscher1996formulation, pototschnig_implementation_2021, Yuwono2024, Chamoli2025, Brandejs2025 offers a flexible framework for obtaining accurate energies and properties which treats electron correlation and relativistic effects on the same footing. It can thereby be used to investigate systems of interest across the periodic table, and benchmark more approximate methods.
In recent years there have been extensive efforts by different groups to address states beyond closed-shell ground states through EOMCC theory for excitation and ionization energies, and electron affinitiesPathak:2016ck, Pathak:2015he, Pathak:2014gv, Pathak:2014dm, Pathak:2016dk, Klein:2008hx, Yang:2012gd, Wang:2015jw, Epifanovsky:2015hsa, Cao:2016fx, Cao:2017id, Zhang:2017jj, Akinaga:2017hv, Wang:2016hx, shee2018equation, Koulias2019, Halbert2021, Halbert2023, Zhang2024, Koulias2019, Zhang2024b, Uhlov2024, Yuwono2025, Li2025, Banerjee2025, Mukhopadhyay2025. While some of these implementations allow for the calculation of transition moments or static properties (through the application of finite fields or via approximate expectation values), full-fledged implementations of CC or EOMCC response theory are much more scarce. Yuan_Xiang_LR have reported an implementation capable of carrying out frequency-dependent LR-CC and LR-EOMCC calculations for different (from 1- to 4-component based) Hamiltonians for electric, magnetic and mixed electric-magnetic properties. A comparison of CC and EOMCC showed that for dipole polarizabilities the two approaches do yield similar results as relativistic effects become more pronounced, but for indirect spin-spin coupling constants there is growing difference between the two. Chakraborty2025 reported an efficient implementation of LR-CC for the spin-free eXact 2-Component Hamiltonian (X2C) and applied it to investigate the static and dynamic dipole polarizabilities of an extensive set of molecular systems.
Beyond linear response, the analytical treatment of quadratic response coupled cluster theory would be particularly useful in applications in which finite field calculations are not straightforward, easy to converge or inapplicable, such for frequency-dependent calculations of non-linear optical properties or in the study of birefringences. However, general implementations beyond linear response are scarcer still, with our report of relativistic QR-EOMCC Yuan_Xiang_QR being, to the best of our knowledge, the first to appear in the literature. This implementation allowed us to gauge the performance of more approximate approaches such as TDDFT and TDHF in the calculation of hyperpolarizabilies, two-photon absorption cross-sections and Verdet constants. However, as discussed above QR-CC and QR-EOMCC are not equivalent, and it is desirable to further characterize the behavior of the latter with respect to the former across the periodic table, before carrying out more extensive applications with the more approximate QR-EOMCC scheme.
The aim of the current work is therefore to report our implementation of quadratic response coupled cluster for relativistic Hamiltonians, capable of handling for both static and frequency-dependent perturbations involving electric or magnetic operators. We showcase it by providing a first comparison of how QR-CC and QR-EOMCC behave as one goes towards the bottom of the periodic table. We thereby chose to re-investigate the first dipole hyperpolarizabilities of hydrogen halides, and the Verdet constants of heavy noble gas atoms as they show significant relativistic effects Yuan_Xiang_QR and include an exploration of the Verdet constants of selected hydrogen halides. Going further than this previous work, we carried out a more extensive exploration on the effect of the Hamiltonian on these properties, and investigated the properties of systems made up of superheavy elements, such as the Tennessine hydride molecule and the Oganesson atom.
Tennessine and Oganesson, whose existence has been recognized by IUPAC and IUPAP in 2015, are the two heaviest elements known. They have been synthesized at the Joint Institute for Nuclear Research in Dubna by nuclear fusion upon bombardment of Berkelium and Californium targets by between atoms of Calcium-48 ions, respectivelyKarol2016a, Karol2016b. While they are expected to be members of groups 17 and 18, the difficulties in their production–only a few atoms are generated at a time, and their lifetimes are in the order of milliseconds–present formidable challenges to the experimental characterization of their physico-chemical properties Smits2023b, making theoretical work an indispensable tool to learn about their electronic structure Smits2023. In view of their large polarizabilties and more pronounced spin-orbit coupling than that in 6p, 5d and 5f block elements, they are also of potential interest as probes for any subtle difference between CC and EOMCC in the calculation of response properties.
2 Theory and Implementation
2.1 Quadratic Response Theory
As was the case in our previous work on coupled cluster linear responseYuan_Xiang_LR (LR-CC) and EOM-CC quadratic response Yuan_Xiang_QR (QR-EOMCC), our work is carried out within the framework of the quasienergy formulation of response theory Christiansen1998. For the sake of brevity, we only review the formalism for CC quadratic response (QR-CC), and refer to the SI for the working equations of QR-EOMCC as presented in Yuan_Xiang_QR. Throughout the text, the CC acronym will imply the use of coupled cluster singles and doubles (CCSD) wavefunctions.
We shall remain within the no-pair approximation visscher1996formulation, liu_going_2013, Almoukhalalati2016, with the electronic part of the unperturbed molecular Hamiltonian given by
where and denote general (occupied or positive energy virtual) indexes, denotes matrix elements of the one-body operator , and the anti-symmetrized matrix elements of the two-body operator . In the calculations reported in this work, corresponds to Coulomb operator. In our current implementation, Gaunt or Breit interactions are assumed to be treated in an effective mannerSchimmelpfennig.Gropen.1998f8, sikkema2009molecular, iliavs2007infinite, saue2002four. External time-(in)dependent perturbations will be represented by one-body operators
where .
Due to the fact that in CC response theory the time-dependent quasienergy Lagrangian Christiansen1998 is made stationary with respect to the wavefunction parametrization (CC amplitudes and multipliers), we can use the () and () rules in perturbation theoryChristiansen1998, helgaker_molecular_2014. The calculation of a third-order property such as quadratic response, thus only requires the determination of the unperturbed and first-order perturbed CC amplitudes and Lagrange multipliers.
The unperturbed CC amplitudes and Lagrange multipliers are defined as
with for particle lines and for hole lines. As a shorthand notation, in the following we shall use greek letters to represent the strings of second quantization operators, with subscripts to represent single or double excitation; for instance, whereas will represent a generic (either single or double) excitation operator. Conversely, will denote a de-excitation operator, which is biorthogonal to operator , that is, , with representing the reference state, here a single Slater determinant made up of molecular orbitals–or molecular spinors whenever the Hamiltonian accounts for spin-orbit coupling. Also, will denote an excited determinant () contained within the singles and doubles manifold.
From the exponential ansatz for the CC wavefunction shavitt2009many, bartlett2007coupled
the time-independent Dirac (Schrödinger) wave equation
and the operator
we have the unperturbed amplitudes and multipliers obtained from the equations
where is matrix representation of the normal-ordered similarity transformed Hamiltonian. The definitions of and are given in Table 1. Further details on the evaluation of these equations in our implementation can be found in pototschnig_implementation_2021.
After solving for the unperturbed wave function parameters, the first-order parameters are next obtained by solving the response equations Christiansen1998, Yuan_Xiang_LR, Yuan_Xiang_QR
Iyt^_, y¯t^_^Yω_YF,ξ^Y, η^Yω_Z = -(ω_X + ω_Y)P( X(ω_X), Y(ω_Y), Z(ω_Z))f^XYZ(ω_X,ω_Y,ω_Z)O^X,YO^¯X,YC^±ω P ¯HFF^X,A^Y,GBG xt^_ yt^_ zt^_F^X,A^YB
2.2 Implementation details
Our QR-CC implementation has been carried out in the ExaCorr module pototschnig_implementation_2021 of the relativistic quantum chemistry program DIRACsaue2020dirac. Below we provide details on the implementation of the QR-CC response function calculation, and refer the reader to prior works concerning (a) the solution of ground-state amplitudes and multipliers pototschnig_implementation_2021; and (b) the solution of first-order response equations Yuan_Xiang_LR, which leveraged our prior implementation of a generalized Davidson eigenvalue solverhirao1982generalization, shee2018equation to implement the solution of systems of linear equations olsen1988solution. We recall that through the use of the TAL-SH tensor library githubRepo2023, a modular component of the ExaTENSOR library lyakh_domainspecific_2019, all tensor operations in our code are parallelized via OpenMP and can be offloaded to NVIDIA or AMD GPUs whenever these are available.
In our implementation, we divide the calculation into different levels of subroutines: (a) a top-level driver to organize calculations for and and to apply the symmetrization operator as in Equation 10k; (b) a driver that calculates for all operators under consideration, solves all corresponding linear response equations for either QR-CC and QR-EOMCC, and calls the driver to calculate the appropriate response function.
In the case of QR-CC, the response function driver carries out the six permutations between the triplet of operators; it calls a subroutine that calculates the terms within the curly braces in subsection 2.1 for a given pair or operators and stores it on set of 2-/4-dimension tensors (singles and doubles manifolds) and performs the contraction with the perturbed amplitudes. This allows us to, if desired, extract the contributions from each of the four terms (for each triplet of operators) and further break these down in terms of the contributions from the singles and doubles manifolds.
The correctness of our QR-CC implementation was validated through a comparison of non-relativistic results (for the HF, HCl and HBr systems) with (a) the analytic implementation of QR-CC in the Dalton aidas2014d, Dalton2020p1 code (for , see LABEL:SI-tab:verification-analytic-hf, LABEL:SI-tab:verification-analytic-hcl and LABEL:SI-tab:verification-analytic-hbr); and (b) finite-field calculations with the ExaCorr code, where the z-component of the electric dipole was employed as perturbing operator (with different perturbation strenghts) and included to the coupled cluster energy calculations. With that, we ensure that the numerical derivatives values (for , see LABEL:SI-tab:compare-nrdz-hf-hbr) are consistent with the orbital-unrelaxed formulation of CC response. We also employed the analytic implementation of LR-CC in the CFOUR code matthews2020coupled to provide an additional check for , and to investigate the relative importance of orbital relaxation for these systems (at the time of writing the QR-CC functionality in CFOUR is not part of the public release). In these calculations, carried out for HF, HCl and HBr, we employed contracted Dunning basis sets, a point-charge nucleus, and correlated all electrons and virtual orbitals. We note that we have at the same time validated the correctness of the code to calculate the left and right-hand side quadratic response equations, by calculating the components of the dipole polarizability with the asymmetric formulation of linear response theory (see e.g. Yuan_Xiang_LR) using the perturbed amplitudes and multipliers that go in the calculation of the QR-CC response functions. The results and computational procedures associated to these validations are found in the SI.
Our code currently lacks the functionality to obtain QR-EOMCC molecular properties from finite-field calculations with the EOMCC model, and implementing such functionality is beyond the scope of this work. As explained in Yuan_Xiang_QR, we had verified our QR-EOMCC implementation against static values reported by coriani_molecular_2016, obtained with a QR-EOMCC implementation in the Dalton code (which, at the time of writing, is not publicly available). For convenience, a comparison between codes for the HF molecule is presented in the SI.
3 Computational details
All QR-CC, QR-EOMCC calculations were carried out with a development version (revision number 8ef365b7) of the DIRAC codesaue2020dirac, DIRAC25. All coupled cluster calculations have been carried out in GPU-accelerated nodes.
As done in our prior EOM-QR calculations Yuan_Xiang_QR, we have used some of the functionality connected to the Cholesky-decomposition approach in ExaCorrpototschnig_implementation_2021 to avoid the explicit construction of the 2-electron repulsion integral tensor in AO basis in the integral transformation step and with that reduce the memory footprint of our calculations. We employed a threshold of 10-5 for all systems except HAt and HTs, for which the threshold was set to 10-4. The outcome of response calculations is not significantly affected by these approximations.
In ref. Yuan_Xiang_QR, which describes our QR-EOMCC implementation, we investigated for the HF molecule the effect of basis set quality and virtual space truncation on the hyperpolarizabilities, and arrived at the conclusion that with a setup consisting of triply augmented triple-zeta basis sets (the Dunning basis set kendall1992electron, woon1993gaussian, wilson1999gaussian for H and the Dyall basis sets dyall2006relativistic, basis-Dyall-TCA2012-131-3962 for all other atoms) was sufficiently converged for investigating the trends across the periodic table. In some cases and for elements up to bromide the t-aug-cc-pVTZ set was used to make a direct comparison to our prior EOM-QR calculations Yuan_Xiang_QR possible. The diffuse functions in question were automatically generated by DIRAC as an even-tempered sequence based on the most diffuse exponents for all angular momenta primitives in the original basis. In this manuscript we provide additional elements to show this choice of basis set was justified by carrying out a basis set study for the HF, HCl and HBr molecules.
In addition to the non-relativistic Levy-Leblond Hamiltonian (LL) levy1967nonrelativistic, visscher2000approximate(as activated by the .Levy-Leblond keyword) and eXact 2-Component (X2C) relativisticiliavs2007infinite, in the coupled cluster calculations we also investigated the molecular mean-field X2C method based on the Dirac-Coulomb (2DCM) and Dirac-Coulomb-Gaunt (2DCGM) Hamiltonians, as well as the spin-free (SF) variants of the X2C and 2DCGM Hamiltonians. In what follows, we shall use the term orbital as shorthand for both spinors and spin-orbitals, depending on the Hamiltonian used in the calculation. In X2C calculations including spin-orbit coupling, we make use of atomic-mean field SOC integrals, obtained from a calculation employing the Douglas-Kroll-Hess Hamiltonian.
For the correlation spaces, these included the valence electrons (for a total 8 electrons corresponding to those from the 1s shell of hydrogen, the ns2np5 shells for the halides and the ns2np6 shells for noble gases) and virtual orbitals with energies up to 5 a.u., a value that was also established in ref. Yuan_Xiang_QR on the basis of calculations for HF. To illustrate the effect of this choice on the results, For HF, HCl and HBr we have also explored larger virtual cutoffs (10 and 100 a.u.) and no cutoffs (all virtuals correlated), and carried out calculations correlating all electrons and all virtual orbitals employing the 2DCM Hamiltonian. Additionally, in the case of the noble gases (Xe, Rn, Og) we also explored the correlation of the d shell.
The relativistic and non-relativistic calculations have been carried out with the Gaussian typevisscher1997dirac and point charge nucleus model, respectively.
The molecular structures employed in our calculations have been taken from huber1979constants for HX (X=F, Cl, Br, I), from pereira_gomes_influence_2004 for HAt, and from deFarias2017 for HTs, so respectively: H-F (0.91680 Å), H-Cl (1.27455 Å), H-Br (1.41443 Å), H-I (1.60916 Å), H–At (1.722 Å) and H-Ts (1.9676 Å). As the choice of orientation in these systems may affect the signs of the electric dipole and first hyperpolarizability (but not their absolute value), we note our results have been tabulated following the convention that the heavy atoms are placed at the origin of the coordinate system, with the hydrogen atom placed on the positive z-axis of the coordinate system.
EOM-CC excitation energy calculations exploiting point-group symmetry were carried out for the noble gas atoms using the RELCCSD module of DIRAC with the 2DCM and the same basis sets and correlating spaces as employed for the QR-CC calculations.
All inputs, outputs and structures used in this paper are available in the Zenodo repository as supplemental information.
4 Results and discussion
In the sections below we discuss our results for the first hyperpolarizability and Verdet constant for systems across the periodic table. We note that the choice of computational setup concerning the basis sets and correlation space, as well as a comparison to the B3LYP functional, have been discussed in detail in our work discussing the implementation of QR-EOMCC Yuan_Xiang_QR. Therefore, in the following we will mostly focus on a comparison between QR-CC and QR-EOMCC for different Hamiltonians.
4.1 First hyperpolarizability of hydrogen halides (HX, X=F–Ts)
Due to the symmetry of the systems under consideration, we restrict ourselves to the calculation of the parallel component of the static first hyperpolarizability ()shelton_measurements_1994
where the element is given by .
In the following we only calculated two unique components, and , so that and are calculated from
respectively. Our results for the frequency-independent case are summarized in Table 2 for and in Table 3 for . To illustrate the trends for the individual components, in Figure 1 we present results for the SF-2DCM and 2DCM Hamiltonians, while in LABEL:SI-tab:hyperpolar-table1 and LABEL:SI-tab:hyperpolar-table2 we report the values for the and components respectively for all Hamiltonians under consideration.
We recall that our use of a Gaussian nuclear model and Dyall bases in all calculations makes for a slightly different setup than that of Yuan_Xiang_QR. For ease of comparison, we also provide results obtained with the Dunning basis set on the halogens, and note small discrepancies for calculations with the Levy-Leblond Hamitonian arising from the difference in nuclear model.
Focusing first on the QR-EOMCC and QR-CC results for the different Hamiltonians for both and , we observe that for HF and HCl scalar relativistic effects are dominant, with spin-orbit effects representing less than 1% of the total response, something to be expected for these molecules which contain only light elements. For HBr we see that spin-orbit coupling is of some importance–in absolute terms, the hyperpolarizabilities increase with respect to spin-free calculations by about 5% (20%) for QR-CC (QR-EOMCC) and reach about 8% for HI and 10% for HAt. For HTs, on the other hand, spin-orbit coupling is the dominant effect, with nearly a twofold increase in value compared to spin-free calculations. For HF to HI the absolute values for are typically rather small, and become larger for HAt and HTs.
| Method | Hamiltonian | HF | HCl | HBr | HI | HAt | HTs |
| CC | LL | -8.0441 | -8.2236 | -5.5096 | 1.5590 | 9.9337 | 46.4544 |
| LL† | -8.0370 | -8.7902 | -5.8960 | nc | nc | nc | |
| SF-X2C | -8.0483 | -8.0824 | -4.1250 | 6.3168 | 35.6445 | 173.6980 | |
| SF-2DCM | -8.0483 | -8.0821 | -4.1253 | 6.3164 | 35.6457 | 173.9446 | |
| X2C | -8.0537 | -8.1031 | -4.3473 | 5.7901 | 39.5212 | 476.3573 | |
| X2C† | -8.0407 | -8.6627 | -4.7106 | nc | nc | nc | |
| 2DCM | -8.0498 | -8.0942 | -4.3343 | 5.7341 | 39.7956 | 495.1556 | |
| 2DCM(1) | -7.9730 | -8.1041 | -4.3898 | nc | nc | nc | |
| 2DCM(2) | -7.9327 | -8.0703 | -4.3439 | nc | nc | nc | |
| 2DCM(3) | -7.9328 | -8.0703 | -4.3441 | nc | nc | nc | |
| 2DCM,∗ | -7.9142 | -8.0537 | -4.7243 | nc | nc | nc | |
| 2DCGM | -8.0494 | -8.0918 | -4.3212 | 5.7574 | 39.8091 | 493.6418 | |
| EOM | LL | -7.4972 | -5.6034 | -0.2535 | 11.1413 | 23.5288 | 66.9526 |
| LL† | -7.5315 | -6.2493 | -0.6946 | nc | nc | nc | |
| SF-X2C | -7.4993 | -5.4499 | 1.1741 | 15.6862 | 46.4367 | 164.0866 | |
| SF-2DCM | -7.4994 | -5.4502 | 1.1731 | 15.6839 | 46.4288 | 164.1359 | |
| X2C | -7.5046 | -5.4712 | 0.9285 | 14.8533 | 45.5480 | 405.8769 | |
| X2C† | -7.5333 | -6.1101 | 0.5181 | nc | nc | nc | |
| 2DCM | -7.5009 | -5.4627 | 0.9445 | 14.8124 | 45.8277 | 423.0797 | |
| 2DCM(1) | -7.6062 | -5.5944 | 0.7430 | nc | nc | nc | |
| 2DCM(2) | -7.6026 | -5.5532 | 0.7973 | nc | nc | nc | |
| 2DCM(3) | -7.6025 | -5.5534 | 0.7966 | nc | nc | nc | |
| 2DCM,∗ | -7.6050 | -5.6405 | -0.4133 | nc | nc | nc | |
| 2DCGM | -7.5016 | -5.4658 | 0.9341 | 14.7813 | 45.8455 | 422.7453 | |
| LL | -0.5469 | -2.6203 | -5.2561 | -9.5822 | -13.5951 | -20.4982 | |
| LL† | -0.5055 | -2.5409 | -5.2014 | nc | nc | nc | |
| SF-X2C | -0.5490 | -2.6324 | -5.2991 | -9.3694 | -10.7921 | 9.6114 | |
| SF-2DCM | -0.5490 | -2.6319 | -5.2985 | -9.3675 | -10.7831 | 9.8088 | |
| X2C | -0.5491 | -2.6319 | -5.2758 | -9.0631 | -6.0268 | 70.4803 | |
| X2C† | -0.5074 | -2.5526 | -5.2287 | nc | nc | nc | |
| 2DCM | -0.5490 | -2.6316 | -5.2787 | -9.0782 | -6.0321 | 72.0758 | |
| 2DCM(1) | -0.3668 | -2.5097 | -5.1328 | nc | nc | nc | |
| 2DCM(2) | -0.3301 | -2.5171 | -5.1412 | nc | nc | nc | |
| 2DCM(3) | -0.3303 | -2.5169 | -5.1407 | nc | nc | nc | |
| 2DCM,∗ | -0.3092 | -2.4132 | -4.3109 | nc | nc | nc | |
| 2DCGM | -0.5478 | -2.6261 | -5.2553 | -9.0239 | -6.0364 | 70.8965 |
In all cases, there are no significant differences between the (SF-)X2C and (SF-)2DCM Hamiltonians, in spite of the more approximate starting point of the X2C calculations, and the effect of the Gaunt interaction is only non-negligible for the HAt and HTs molecules (though it only contributes to about 1%).
We have noticed small differences between our results and those previously reported Yuan_Xiang_QR, the latter showing larger (in absolute value) hyperpolarizabilties. We have traced this difference to the use of Dyall basis sets for the current calculations, as opposed to the Dunning basis sets. This difference can be understood as a consequence of the bias towards somewhat less diffuse exponents as part of the process of obtaining relativistic energy-optimized basis sets.
Focusing now first on QR-CC values for , for HF to HBr we observe negative values (similar in magnitude for HF and HCl, but less negative for HBr), and positive for HI, HAt, and HTs, with rather large values for the latter molecule. These trends result from the fact that the clearly negative component for HF initially becomes less negative as the halogen becomes heavier, changes sign for HAt, and becomes strongly positive for HTs. The component shows a similar trend, with positive values from HBr onwards, but is smaller in magnitude than for most systems. These trends are qualitatively the same for QR-EOMCC across the series, with as important difference that the sign change in upon going down the series already occurs for HBr.
Quantifying the differences between QR-CC and QR-EOMCC by computing (), we see that differences are small for HF but become non-negligible already for HCl and remain sizable for the other species. In relative terms, the differences for are about 6% for HF, but grow to about 30% for HCl, 120 % for HBr and 150% for HI. For the very large positive hyperpolarizability of HTs, the difference is in absolute sense the largest, but amounts to only 15% of the 2DCGM value. For the more accessible HAt molecule, the effect is similar but less pronounced. We note the situation is a bit different for , where we observe that apart from HTs, the difference between is QR-CC and QR-EOMCC is fairly modest across the series.
The results above have been obtained with a computational setup (triply augmented basis sets and virtual spaces truncated at 5 a.u.) established on the basis of a verification of basis set and correlation effects for the HF molecule in our paper describing the implementation of QR-EOMCC Yuan_Xiang_QR, and aimed at consistently addressing the whole hydrogen halide series. It is nevertheless important to have an idea of how close this setup is to the converged basis set and correlation treatment. To this end, we have investigated the effect of varying the number of diffuse functions and the cardinal number of the basis sets for HF, HCl and HBr, employing the 2DCM Hamiltonian.
From our results, found in LABEL:SI-tab:InfluenceOfDiffuseSI, we have that for triple-zeta basis sets, the triply and quadruply augmented basis sets typically show agreement better than 0.1 a.u. for all molecules, effectively validating our choice and in line with prior work on the light hydrogen halides Fernandez1998. We consider nevertheless that doubly augmented triple-zeta basis set can still be a good choice whenever it is not possible to employ larger basis sets for heavier elements. Furthermore, as expected singly augmented triple zeta bases are not sufficiently flexible, but are definitely much better suited than any double-zeta basis.
Another surprising but very interesting behavior revealed by this benchmark is that QR-CC and QR-EOMCC results are actually very close to each other for double-zeta basis sets without diffuse functions–including for HCl and HBr–and as soon as diffuse functions are added the differences between the two methods start to grow. To the best of our knowledge, this point has not been explored in the literature in detail, and possibly is the reason why our results seem at first at odds with the claims in the literature that CC and EOMCC yield results which are typically very close to each other.
With respect to the convergence of the correlation space, we observe from Table 2 and Table 3 that the 5 a.u. threshold is not a poor choice, but it is clear that with it the virtual space is not converged. Upon exploring other truncations while correlating only the valence electrons, we see that at 10 a.u. results are not quite converged either, whereas with a truncation at 100 a.u. results are essentially equivalent to those correlating all virtuals. An important finding is that, in qualitative terms, there is no difference in the trends across the halogen series for all virtual truncations employed. Finally, for HBr core correlation is not negiglible, accounting for an increase of 4% and 18% for calculated with QR-CC and QR-EOMCC. This change reduces the difference between the two approaches by nearly 1 a.u., but in any case the discrepancy remains significant (larger than 4 a.u.). It is important to keep in mind that such change is driven primarily by an increase in .
Going beyond the static case, in Figure 2 we present first the results for second harmonic generation (SHG) calculations on HI, which corresponds to the case in which . As this property has been discussed in detail in our work presenting QR-EOMCC Yuan_Xiang_QR, we refer the reader to it for further details on the behavior of the dispersion curves. We nevertheless highlight the fact that that in the region around a.u., corresponding to half the excitation energy of the spin–orbit allowed transition towards the state, we see the expected pole for the quadratic response functions. Beside that, we observe that across the range of frequencies sampled, the difference between QR-CC and QR-EOMCC results amounts to a nearly constant shift for low frequencies. For the region close to the resonance, on the other hand, we see a relatively small decrease in the difference between QR-CC and QR-EOMCC.
In Table 4 we present a second dynamic property, the electro-optical Pockels effect boyd2020nonlinear, norman_principles_2018, which corresponds to the condition . The Pockels effect can be understood as capturing the change in the dynamic polarizability due to an applied static electric field. A first consequence of the condition is that the resonances now only occur near the first excited states (e.g. for HI, at about a.u.) instead of at smaller frequencies such as in SHG. This means that for the systems under study we are very much off-resonance and consequently the values remain very similar to those in the static case. Interestingly, we see a (very small) decrease in the difference between QR-CC and QR-EOMCC as the perturbation frequency increases, as found for the SHG case, but that in no way is capable of reversing the growing difference between QR-CC and QR-EOMCC across the halide series, and that these differences are driven by the component.
| Method | Hamiltonian | HF | HCl | HBr | HI | HAt | HTs |
| CC | LL | -4.9459 | -10.3349 | -14.5008 | -16.3776 | -18.5536 | -50.1339 |
| LL† | -4.6309 | -10.0969 | -14.3693 | nc | nc | nc | |
| SF-X2C | -4.9455 | -10.4108 | -15.4252 | -20.1042 | -41.2161 | -191.5107 | |
| SF-2DCM | -4.9454 | -10.4091 | -15.4184 | -20.0927 | -41.1858 | -191.6326 | |
| X2C | -4.9536 | -10.4424 | -16.0964 | -24.8101 | -83.2791 | -599.5770 | |
| X2C† | -4.6278 | -10.1948 | -15.9223 | nc | nc | nc | |
| 2DCM | -4.9467 | -10.4348 | -16.0811 | -25.0659 | -84.7735 | -618.4277 | |
| 2DCM(1) | -5.1487 | -10.5333 | -16.2143 | nc | nc | nc | |
| 2DCM(2) | -5.1251 | -10.5378 | -16.2212 | nc | nc | nc | |
| 2DCM(3) | -5.1252 | -10.5381 | -16.2215 | nc | nc | nc | |
| 2DCM,∗ | -5.1300 | -10.6165 | -16.5150 | nc | nc | nc | |
| 2DCGM | -4.9478 | -10.4450 | -16.1219 | -25.1406 | -84.9810 | -617.7646 | |
| EOM | LL | -5.2227 | -10.2724 | -14.6625 | -15.0853 | -16.4412 | -43.4775 |
| LL† | -4.9146 | -9.9864 | -14.4596 | nc | nc | nc | |
| SF-X2C | -5.2236 | -10.3699 | -15.8135 | -19.9561 | -43.5784 | -195.9931 | |
| SF-2DCM | -5.2234 | -10.3687 | -15.8100 | -19.9571 | -43.5978 | -196.3284 | |
| X2C | -5.2315 | -10.4017 | -16.4911 | -24.6949 | -85.2353 | -553.0503 | |
| X2C† | -4.9126 | -10.1052 | -16.2428 | nc | nc | nc | |
| 2DCM | -5.2248 | -10.3948 | -16.4807 | -24.9707 | -86.8113 | -570.8098 | |
| 2DCM(1) | -5.3962 | -10.4594 | -16.5684 | nc | nc | nc | |
| 2DCM(2) | -5.3543 | -10.4722 | -16.5866 | nc | nc | nc | |
| 2DCM(3) | -5.3544 | -10.4723 | -16.5867 | nc | nc | nc | |
| 2DCM,∗ | -5.3591 | -10.5605 | -16.7843 | nc | nc | nc | |
| 2DCGM | -5.2252 | -10.4008 | -16.5012 | -24.9851 | -86.8673 | -570.0000 | |
| LL | 0.2768 | -0.0624 | 0.1618 | -1.2923 | -2.1124 | -6.6564 | |
| LL† | 0.2838 | -0.1105 | 0.0903 | nc | nc | nc | |
| SF-X2C | 0.2780 | -0.0409 | 0.3883 | -0.1482 | 2.3622 | 4.4824 | |
| SF-2DCM | 0.2781 | -0.0404 | 0.3915 | -0.1357 | 2.4120 | 4.6959 | |
| X2C | 0.2780 | -0.0407 | 0.3947 | -0.1151 | 1.9562 | -46.5267 | |
| X2C† | 0.2848 | -0.0896 | 0.3205 | nc | nc | nc | |
| 2DCM | 0.2781 | -0.0400 | 0.3997 | -0.0952 | 2.0379 | -47.6178 | |
| 2DCM(1) | 0.2475 | -0.0739 | 0.3541 | nc | nc | nc | |
| 2DCM(2) | 0.2292 | -0.0657 | 0.3654 | nc | nc | nc | |
| 2DCM(3) | 0.2292 | -0.0657 | 0.3652 | nc | nc | nc | |
| 2DCM,∗ | 0.2291 | -0.0560 | 0.2693 | nc | nc | nc | |
| 2DCGM | 0.2774 | -0.0442 | 0.3793 | -0.1555 | 1.8864 | -47.7646 |
| System | CC | EOM | % | CC | EOM | % | ||||
| HF | 0.01 | -8.0557 | -7.5067 | -0.5490 | 6.81 | 4.9484 | 5.2266 | -5.4974 | -5.62 | |
| 0.02 | -8.0733 | -7.5242 | -0.5491 | 6.80 | 4.9535 | 5.2320 | -5.5026 | -5.62 | ||
| 0.03 | -8.1028 | -7.5536 | -0.5492 | 6.78 | 4.9619 | 5.2411 | -5.5111 | -5.63 | ||
| 0.05 | -8.1985 | -7.6488 | -0.5496 | 6.70 | 4.9886 | 5.2696 | -5.5382 | -5.63 | ||
| 0.09 | -8.5492 | -7.9982 | -0.5510 | 6.44 | 5.0776 | 5.3652 | -5.6286 | -5.66 | ||
| HCl | 0.01 | -8.1047 | -5.4717 | -2.6329 | 32.49 | 10.4421 | 10.4022 | -13.0750 | 0.38 | |
| 0.02 | -8.1362 | -5.4991 | -2.6371 | 32.41 | 10.4639 | 10.4243 | -13.1009 | 0.38 | ||
| 0.03 | -8.1891 | -5.5452 | -2.6439 | 32.29 | 10.5001 | 10.4610 | -13.1441 | 0.37 | ||
| 0.05 | -8.3626 | -5.6965 | -2.6662 | 31.88 | 10.6151 | 10.5775 | -13.2813 | 0.35 | ||
| 0.09 | -9.0230 | -6.2765 | -2.7465 | 30.44 | 10.9995 | 10.9670 | -13.7460 | 0.30 | ||
| HBr | 0.01 | -4.3414 | 0.9411 | -5.2825 | 121.68 | 16.1002 | 16.5005 | -21.3828 | -2.49 | |
| 0.02 | -4.3629 | 0.9310 | -5.2939 | 121.34 | 16.1579 | 16.5600 | -21.4518 | -2.49 | ||
| 0.03 | -4.3992 | 0.9138 | -5.3130 | 120.77 | 16.2543 | 16.6596 | -21.5673 | -2.49 | ||
| 0.05 | -4.5196 | 0.8553 | -5.3749 | 118.93 | 16.5659 | 16.9815 | -21.9408 | -2.51 | ||
| 0.09 | -5.0011 | 0.6012 | -5.6023 | 112.02 | 17.6886 | 18.1420 | -23.2909 | -2.56 | ||
| HI | 0.01 | 5.7433 | 14.8309 | -9.0875 | -158.23 | 25.1222 | 25.0273 | -34.2098 | 0.38 | |
| 0.02 | 5.7710 | 14.8866 | -9.1155 | -157.95 | 25.2923 | 25.1982 | -34.4079 | 0.37 | ||
| 0.03 | 5.8175 | 14.9800 | -9.1626 | -157.50 | 25.5800 | 25.4873 | -34.7425 | 0.36 | ||
| 0.05 | 5.9679 | 15.2841 | -9.3162 | -156.10 | 26.5371 | 26.4494 | -35.8533 | 0.33 | ||
| 0.09 | 6.5089 | 16.4030 | -9.8941 | -152.01 | 30.4156 | 30.3485 | -40.3097 | 0.22 | ||
In the following we look more closely at the differences between QR-CC and QR-EOMCC by analyzing the contributions from the different terms to the response functions, found in SI (LABEL:SI-tab:breakdown-qrf-cc-beta for QR-CC and LABEL:SI-tab:breakdown-qrf-eomcc-beta for QR-EOMCC respectively), as we go across the series from HF to HI.
For this analysis, it is interesting to also consider the results for the static dipole polarizabilties for the same molecules. The reason for that is twofold: first, it is well-established that the formal difference between LR-CC and LR-EOMCC is connected to the presence of the term in the LR-CC response function (and therefore the right-hand side ground to excited state transition matrix elements), whereas LR-EOMCC employs a simpler expression–for both methods perturbed amplitudes are obtained from eq. Equation 10a, and left-hand side transition matrix elements are formally the same. This means that a large difference in polarizability is indicative of the relative importance of the full term, and consequently of an important difference in the right-hand side ground-to-excited state transition matrix elements.
Second, if one considers a sum-over-states expression for the hyperpolarizabilities, we have in the denominator one has electronic state energies (which are identical for CC and EOMCC), while in the numerator appears the product of ground-to-excited state transition matrix elements (from linear response) and state to state transition matrix elements ( LABEL:SI-state-to-state-tm, given in terms of the contractions found in subsection 2.1). The consequence of the expression of the numerator is that, if the linear response terms for CC and EOMCC are similar while the full quadratic response function is not, the difference should be coming to the state to state transition matrix elements; or, in other words, in the description of transition moments between excited states.
The dipole polarizabilities, shown in Figure 3 (values for dipole moments and polarizabilities are reported in LABEL:SI-tab:polar-hx-table) increase across the series, which is consistent with the increase in size of the halides accompanied by their decreasing electronegativity. The differences between CC and EOMCC do increase across the series for the and components and consequently, for the mean polarizability as well. However, in contrast to the hyperpolarizabilities, in relative terms these differences are roughly the same and of about 2% for and 1.5% for .
The difference between CC and EOMCC is more strongly felt for the anisotropy, which is about 3% for HF, but 8-12% for HCl to HAt, and reaches about 30% for HTs, though in absolute terms the differences are still modest. An intriguing observation is that the polarizability anisotropy, while becoming less important for the almost spherical HAt and HTs molecules, changes sign upon going from HAt to HTs. This means that HTs can be more easily polarized in the direction perpendicular to the bond while the reverse is true for the lighter homologues that are more easily polarized along the bond. The sign change occurs with spin-orbit coupling but not without (though in magnitude the spin-orbit and spin-free anisotropies are very close to each other), making further studies necessary to better characterize it. We note that our CCSD dipole moments compare well with available experimental NISTconstants, Hohm2013 or theoretical work (for HTs) thierfelder_scalar_2009, though for reaching benchmark accuracy we would need to account for higher excitations, something which is also beyond the scope of this work. These findings add to the body of data supporting the argument that differences between LR-CC and LR-EOMCC properties are in practice small, and in particular when correlating the same number of electrons.
As mentioned above, another consequence of the similarity between CC and EOMCC polarizabilities is that it suggests their difference arises from the state to state transition matrix elements. While at this time we have not yet implemented the calculation of excited state properties (and such implementation is beyond the scope of this work), we can nevertheless gain some insight by considering the individual contributors to the value of the QR-CC functions.
We see that across the halide series the dominant contributions typically arise from the terms from eq. subsection 2.1, involving both the perturbed amplitudes and multipliers ( and ). Furthermore, we observe that these two contributions to the final value in general tend to counteract each other, and that the magnitude of the “doubles” contributions increases as the halide becomes heavier. Terms from eq. subsection 2.1, involving only perturbed amplitudes and ground-state multipliers ( and ) show smaller (but non-negligible) net contributions than those above. They also increase in absolute value as the halide becomes heavier.
The relative importance of contribution of the terms in eq. subsection 2.1 also changes depending on which response function we look at. For , the term is the one which contributes the most, but for there are more important cancellations between the “singles” and “doubles” contributions from it so that the term becomes a more important net contributor, being the dominant one for the HI molecule followed by , whereas and show similar contributions.
For QR-EOMCC, we see some similarities to the QR-CC picture, in that contributions beyond “singles” become increasingly important as the halogen becomes heavier, while the term is typically a key contribution and dominates for the lightest elements. The value of the contribution also appears to follow closely that of in QR-CC. The terms and , on the other hand, become increasingly more important for heavier systems, and for of HBr and HI, they offset the other contribution and are largely responsible for the sign changes between QR-CC and QR-EOMCC for these systems.
4.2 Verdet constant of noble gases (Xe-Og) and hydrogen halides (HF-HI)
In Table 5 we present our results for the calculation of the Verdet constant for Xe, Rn and Og. In line with our prior work, we considered three laser wavelengths (589.0 nm, 694.3 nm, and 1064.0 nm) to compare QR-EOMCC and QR-CC for different Hamiltonians. The Verdet constants are evaluated as quadratic response functionscoriani_coupled_1997, norman_quadratic_2004:
with the number density of the gas, the elementary charge, the electron mass, the speed of light in vacuo, and is magnetic dipole moment operator. For the noble gases, due to atomic symmetry, we only needed to calculate one unique component and could obtain through the expression
Comparing first the differences between QR-CC and QR-EOMCC, we observe that these become larger as the wavelength decreases, and increase as the atoms become heavier. However, if we consider the magnitude of these differences, we see that the differences are rather small overall (about 1 to 1.4%). Furthermore, the differences are very similar for the different wavelengths. These additional calculations with respect to those already reported by Yuan_Xiang_QR, allow us to more clearly separate contributions from different relativistic effects. With respect to the Hamiltonian and as discussed previously, the Verdet constant for Xe is still well described by non-relativistic Hamiltonians cadene_circular_2015, Yuan_Xiang_QR, and scalar relativity is the dominant relativistic effect. Furthermore, we observe no difference between (SF-)X2C and (SF-)2DCM results, and find only rather small differences X2C and 2DCM. As this is an atomic system, these differences are primarily due to the difference in Hamiltonian used to compute the spin-orbit screening in the mean-field approximation used in the X2C approach. Adding the Gaunt interaction to this mean-field screening acts to slightly decrease the 2DCM value, in the direction of the spinfree and X2C (Douglas-Kroll-Hess based screening) result. We also observe that in spite of the fact that QR-CC values are smaller than QR-EOMCC ones, they still overestimate the experimental results. For Rn relativistic effects are important and spin-orbit coupling effects significantly increase the Verdet constant. We again see some small differences between X2C and 2DCM. For Og, spin-free results with X2C and 2DCM are roughly the same, indicating that the bare nucleus and screened X2C transformation matrices are very similar (this is the only difference between the two spin-free approaches, see sikkema2009molecular for a more extensive discussion). With spin-orbit coupling differences between X2C and 2DCM are more noticeable at about 3%. For this superheavy element, the spin-orbit effect is very significant and even dominates over scalar relativity, with 2DCM values being roughly three times that of SF-2DCM. While much larger in magnitude, the Gaunt interaction behaves similarly in Og as in Xe, reducing the Verdet constant at all frequencies.
In Figure 4 and LABEL:SI-tab:verdet-table-scan-nsnp we present the frequency dependence of the Verdet constant obtained with the 2DCM Hamiltonian for additional frequencies than those in Table 5. From these data, we observe that no significant discrepancy arises from the use of QR-EOMCC instead of QR-CC. Across the range of studied wave lengths, the differences remain on the order of 1.1-1.3%.
The magnitude of the Verdet constant for Rn is found to be roughly twice of that for Xe across the range of studied wavelenghts, while the one of Og is much larger still, with its magnitude being 4.5 times larger for the longest to about 6.5 times larger for the shortest wavelenghts. This is in stark contrast to what is observed for the spin-free calculation presented in Table 5, where we see changes of only a factor of roughly 1.5 going from Xe to Rn, and of 1.8 going from Rn to Og. This is consistent with prior theoretical investigations of Og, that predict this element to be much more polarizable due to the extremely large spin-orbit effects in this superheavy element Jerabek2019a, Jerabek2019b.
While the first excited state is lowered due to relativity, this strongly increased polarizability due to spin-orbit effects manifests itself already far away from this resonance. To document this, we present in LABEL:SI-tab:excitation-ng calculations for the lowest excited states of these atoms. These show that while the first excited state is progressively lowered by spin-orbit coupling as the atom gets heavier, the wavelengths probed are still far away from the resonances.
| (589 nm) | (694.3 nm) | (1064 nm) | ||||||||
| Atom | Hamiltonian | CC | EOM | CC | EOM | CC | EOM | |||
| Xe | SF-X2C | 12.39 | 12.54 | -0.15 | 8.72 | 8.82 | -0.11 | 3.60 | 3.64 | -0.04 |
| SF-2DCM | 12.40 | 12.55 | -0.15 | 8.72 | 8.83 | -0.11 | 3.60 | 3.64 | -0.04 | |
| X2C | 12.95 | 13.10 | -0.15 | 9.10 | 9.20 | -0.11 | 3.74 | 3.79 | -0.05 | |
| 2DCM | 13.00 | 13.16 | -0.15 | 9.14 | 9.24 | -0.11 | 3.76 | 3.81 | -0.05 | |
| 2DCGM | 12.97 | 13.12 | -0.15 | 9.11 | 9.22 | -0.11 | 3.75 | 3.80 | -0.05 | |
| Rn | SF-X2C | 18.79 | 19.01 | -0.22 | 13.14 | 13.30 | -0.15 | 5.38 | 5.44 | -0.06 |
| SF-2DCM | 18.82 | 19.04 | -0.22 | 13.16 | 13.31 | -0.15 | 5.38 | 5.45 | -0.06 | |
| X2C | 25.62 | 25.90 | -0.28 | 17.73 | 17.92 | -0.19 | 7.14 | 7.22 | -0.08 | |
| 2DCM | 26.07 | 26.35 | -0.28 | 18.03 | 18.22 | -0.20 | 7.26 | 7.34 | -0.08 | |
| 2DCGM | 25.95 | 26.24 | -0.28 | 17.95 | 18.15 | -0.20 | 7.23 | 7.31 | -0.08 | |
| Og | SF-X2C | 35.45 | 35.83 | -0.38 | 24.44 | 24.71 | -0.27 | 9.81 | 9.91 | -0.11 |
| SF-2DCM | 35.58 | 35.95 | -0.38 | 24.53 | 24.79 | -0.26 | 9.84 | 9.95 | -0.11 | |
| X2C | 144.66 | 146.52 | -1.85 | 92.02 | 93.20 | -1.18 | 33.34 | 33.77 | -0.43 | |
| 2DCM | 151.00 | 152.95 | -1.95 | 95.74 | 96.98 | -1.24 | 34.55 | 35.00 | -0.45 | |
| 2DCGM | 149.96 | 151.92 | -1.96 | 95.12 | 96.36 | -1.24 | 34.34 | 34.79 | -0.45 | |
To further complete the picture, we computed the 2DCM dipole polarizabilities for the noble gases with LR-CC and LR-EOMCC for the same range of frequencies as for the Verdet constant. The results, presented in LABEL:SI-tab:polar-ng-table, are in line with the findings for the static polarizabilties for the HX systems and our work Yuan_Xiang_LR in that there is no appreciable difference between the two CC variants for this property.
Our results are in good agreement with the experimental static dipole polarizabilities for Xe and Rn (27.3 and 35.8 a.u. respectively) Hohm1994, CRC2014, and likewise for the theoretical predictions for Og (57.98 a.u.) of Smits and coworkers Smits2023. Though we did not calculated values for atomic Ts, our results from HTs capture well the trend of decreasing polarizability from Ts to Og (76.3 to 57.98 a.u.) Smits2023 predicted in that work. We attribute the observed discrepancies to the lack of higher excitations in CCSD.
To check for the contribution of subvalence electrons, we also considered additional calculations in which we retained the virtual space truncation at 5 a.u. but included also the electrons. In our prior QR-EOMCC calculations Yuan_Xiang_QR we found, based on non-relativistic calculations (the difference of calculations with the same correlating space, and by correlating all occupied and virtual orbitals), that the effect of correlating electrons beyond the valence shell for the Xe atom could result in a decrease of the Verdet constant by about -0.08, -0.19 and -0.28 [10-3 rad/(T m)], for 1064 nm, 694.3 nm and 589 nm respectively.
Our current results, shown in LABEL:SI-tab:verdet-table-scan-n-1dnsnp, confirm that a decrease in Verdet constant is indeed observed for all atoms, though with the current virtual space truncation for Xe we recover slightly smaller changes (-0.07, -0,18 and -0.26 [10-3 rad/(T m)], for 1064 nm, 694.3 nm and 589 nm respectively). This is consistent with our earlier finding, that upon correlating the electrons, the excitation energies for all systems are (slightly) shifted upward. Another change seen upon correlating the electrons is that the difference between QR-CC and QR-EOMCC values increases somewhat, though in relative terms these differences remain close to 1% overall.
As a first comparison beyond atomic systems, in Table 6 we present the Verdet constants obtained with the 2DCM Hamiltonian for the hydrogen halides up to HI. The values of the unique response functions used to calculate the Verdet constant are given in the SI. whereas systems.
Due to their axial symmetry, for the HX systems we require at least two response function evaluations; in addition to we calculated (and verified that we obtain that it yields the same result as ), so that the Verdet constants are calculated with the expression
| V() | |||||
| System | CC | EOM | % | ||
| HF | 0.01 | 0.0161 | 0.0162 | -0.0001 | -0.41 |
| 0.02 | 0.0647 | 0.0650 | -0.0003 | -0.41 | |
| 0.03 | 0.1461 | 0.1467 | -0.0006 | -0.41 | |
| 0.05 | 0.4112 | 0.4128 | -0.0017 | -0.40 | |
| 0.09 | 1.3954 | 1.4009 | -0.0056 | -0.40 | |
| HCl | 0.01 | 0.0781 | 0.0789 | -0.0008 | -0.99 |
| 0.02 | 0.3133 | 0.3165 | -0.0031 | -0.99 | |
| 0.03 | 0.7089 | 0.7159 | -0.0070 | -0.99 | |
| 0.05 | 2.0042 | 2.0239 | -0.0197 | -0.98 | |
| 0.09 | 6.9238 | 6.9903 | -0.0665 | -0.96 | |
| HBr | 0.01 | 0.1356 | 0.1370 | -0.0014 | -1.07 |
| 0.02 | 0.5445 | 0.5503 | -0.0058 | -1.06 | |
| 0.03 | 1.2335 | 1.2466 | -0.0131 | -1.06 | |
| 0.05 | 3.5036 | 3.5405 | -0.0369 | -1.05 | |
| 0.09 | 12.3195 | 12.4447 | -0.1252 | -1.02 | |
| HI | 0.01 | 0.2689 | 0.2726 | -0.0037 | -1.37 |
| 0.02 | 1.0813 | 1.0960 | -0.0147 | -1.36 | |
| 0.03 | 2.4545 | 2.4879 | -0.0334 | -1.36 | |
| 0.05 | 7.0169 | 7.1115 | -0.0946 | -1.35 | |
| 0.09 | 25.3084 | 25.6359 | -0.3275 | -1.29 | |
Following the trends observed for the atoms, the Verdet constants increase across the HF to HI series. This is in line with the increase in molecular polarizabilities for the series, and a reflex of the Verdet constant representing the change in dynamic polarizability in the presence of the external static magnetic field. Likewise, we have that QR-CC and QR-EOMCC yield very similar results for a given frequency and molecular system. In relative terms, we see that the differences are on par with those for the noble gases, with the exception of HF which shows even smaller differences. We note that as in the case of SHG and the Pockels effect, we have a small decrease in the difference between QR-CC and QR-EOMCC for larger perturbing frequencies.
In view of the arguments put forward in the analysis of the hyperpolarizabilities, and gven the similarity between CC and EOMCC polarizabilities for these systems–and consequently of their ground to excited transition matrix elements–the good agreement for the the Verdet constants suggests the terms connected to the state to state transition moments ( in subsection 2.1) should also be similar.
In order to assess this hypothesis, we have analyzed the contributions from different contractions to the response functions in the Verdet constant case as well. The breakdown of the contributions are found in LABEL:SI-tab:breakdown-qrf-cc-verdet for QR-CC and in LABEL:SI-tab:breakdown-qrf-eomcc-verdet for QR-EOMCC. From them, we observe that the contributions arise essentially from the term across the series, with an increased contribution from the for QR-CC seen for HI. We observe that contributions to these terms are essentially made up of singles for all species, a behavior which is distinct from the hyperpolarizabilities.
A comparison between the trends across the halide series and those for the Pockels effect is also interesting, as it highlight the difference between the perturbing operators : for Pockels, the driver for the difference between QR-CC and QR-EOMCC is the operator, which drives the response of the polarizability along the molecular axis. For verdet, on the other hand, the magnetic operator induces an action perpendicular to the molecular axis.
To conclude this section, we consider that through the analysis of the quadratic response functions, we have been able to identify the mechanisms through which the differences between QR-CC and QR-EOMCC arise for the systems under consideration. We also show that depending on the property under considreration, QR-EOMCC can offer a very similar quality than QR-CC–and as shown previously, while outperforming density functional approximations such as B3LYP Yuan_Xiang_QR–at a substantially less cost.
5 Conclusions
In this work we have presented an implementation in the GPU-accelerated ExaCorr module of coupled cluster quadratic response (QR-CC) theory which can be used in combination with both (2- and 4-component based) relativistic and non-relativistic Hamiltonians, thus extending our prior implementation quadratic response based on the equation of motion (QR-EOMCC) formalism Yuan_Xiang_QR.
We have showcased this implementation by carrying out a detailed comparison between QR-CC and QR-EOMCC in the calculation of both static and dynamic first hyperpolarizabilities (where we determined dispersion curves for second harmonic generation and the Pockels electro-optical effect) for the hydrogen halide series (for HF to HTs), as well as the calculation of the Verdet constant for noble gases (Xe, Rn and Og) and hydrogen halides (HF to HI). In addition to the comparison of non-relativistic and X2C Hamiltonians as done previously, we have carried out calculations with the X2C molecular mean field approach, and gauged the contribution of spin-orbit coupled and scalar relativistic effects.
For the first hyperpolarizabilities, we find that for the HF molecule the difference between QR-CC and EOM-CC is relatively small (around 4%), in line with the studies reported in the literature employing non-relativistic approaches that do not see significant differences between the two approaches. The differences between QR-CC and QR-EOMCC do however increase as one goes down the periodic table, with differences of about 20% already for HCl and a sign change for in HI. Our findings also point to the fact that such differences are significantly smaller when basis sets without diffuse functions are employed, and are particularly small (from HF to HBr) when double zeta bases such as cc-pVDZ are used. We have attributed this difference to differences in the treatment of terms related to state to state transition moments between QR-CC and QR-EOMCC, based on the analysis of the dominant contributions to the hyperpolarizabilties, and the very good agreement between CC and EOMCC dipole polarizabilities. That said, the overall shape of the dispersion curves remains unchanged, indicating that in qualitative terms the QR-EOMCC model still behaves quite similarly to QR-CC.
The contribution of spin-orbit coupling to the hyperpolarizability is rather small up to HBr, but becomes significant from HI onwards, also for the static case, but the fact that we observe similar trends across the HX series for all Hamiltonians considered indicates that, in spite of the non-additivity between relativisic effects and electron correlation, the differences between QR-CC and QR-EOMCC are driven by electron correlation.
Truncation of the correlating space appears to affect QR-EOMCC more than QR-CC for the component for HBr, but the additional electron correlation recovered does not modify the trends upon going down the series from HF to HBr. We note, however, that results can only be considered converged with respect to virtual space truncation if thresholds are set to higher energies (around 100 a.u. in our tests). Furthermore, core(-valence) as a non-negligible effect on the hyperpolarizability of HBr, calling for more systematic studies. We also found orbital relaxation to be small but non-negligible for the polarizability of HBr (1.7%), and it would be of interest to consider such effect for higher-order properties in future work.
In contrast to the hyperpolarizabilies, for the Verdet constant the difference between QR-CC and EOM-CC is minor and roughly constant percentage-wise (about 1%) for all systems studied. Spin-orbit coupling accounts for a relatively small part of the response (about 4%) in Xe, becomes important for Rn (about 35-40%) and absolutely dominates for Og. For the noble gases, correlating the d electrons in addition to the valence reduces the magnitude of the Verdet constants by about 10%. The comparison of the Verdet constants to the Pockels effect for the hydrogen halides further underscored our argument that the differences between QR-CC and QR-EOMCC are likely driven by the description of state to state transition moments for the operators representing the external static fields (magnetic dipole for Verdet, electric dipole for Pockels). We intend to revisit this point in future work, once our implementation allows for the calculation of excited state properties that may shed further light on these differences.
The calculations presented here, albeit restricted to a relatively small set of atomic and molecular systems serve to highlight similarities, but also sometimes important differences, in behavior between the QR-CC and QR-EOMCC approaches across the periodic table and for different molecular properties. Given the rather favorable computational cost and implementation complexity of QR-EOMCC with respect to QR-CC, it is important to further characterize differences between the two approaches in the calculation of nonlinear properties for heavier element systems.
This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725 (allocations CHM160, CHM191 and CHP109). XY, LH and ASPG acknowledge funding from projects CPER WaveTech, Labex CaPPA (Grant No. ANR-11-LABX-0005-01), ANR CompRIXS (Grant Nos. ANR-19-CE29-0019 and DFG JA 2329/6-1), ANR SCREECHES (Grant Nos. ANR-24-CE29-0904) the I-SITE ULNE project OVERSEE and MESONM International Associated Laboratory (LAI) (Grant No. ANR-16-IDEX-0004), as well support from the French national supercomputing facilities (Grant Nos. DARI A0150801859, A0170801859) and the HPC center at the University of Lille.
The data (input/output) corresponding to the calculations of this paper are available at the Zenodo repositoryzenodo.17215085.
The Supporting Information is available free of charge on the ACS Publications website at DOI: XXX. Working equations for CCSD quadratic response theory; QR-EOMCC equations; Remarks on state to state transition strenghts; Comparison to analytic QR-CC results from other codes; Comparison between analytic QR-CC to finite-field calculations with ExaCorr ; QR-CC and QR-EOMCC and results for different Hamiltonians for HF to HTs systems; Breakdown of the contributions of different terms of the QR-CC and QR-EOMCC response functions for the components, as well as those to obtain the Verdet constant, for HF to HI systems; Basis set convergence study for QR-CC and QR-EOMCC for the components for HF to HBr systems, using the 2DCM Hamiltonian; Dipole moment and polarizabilities for HF to HTs, and Xe to Og systems; Verdet constants for Xe to Og systems; Excitation energies for Xe to Og systems; sample timings for QR-CC and QR-EOMCC response function evaluations.